The ratio of Jack’s allowance to John’s allowance was 1:3. After they each spent $68 on a badminton racket, Jack had 1/4 as much money as John. How much was Jack”s allowance
Answer by greenestamps(13331)



x=Jack's allowance
3x=John's allowance

x-68 = Jack's after he spent $68
3x-68 =John's after he spent $68

Now Jack had 1/4 as much as John:

x-68=(3x-68)/4
4x-272=3x-68
x=204

ANSWER: Jack's allowance was $204

CHECK:
Jack originally: x=$204
John originally: 3x=$612
Jack after spending $68: $136
John after spending $68: $544
$136/$544 = 1/4